You bought a $1,000 par Albany NY 20-year bond with equal annual amortization. How much principal...

  1. You bought a $1,000 par Albany NY 20-year bond with equal annual amortization.

    1. How much principal will you receive each year?

    2. If the coupon rate is 2.8%/year, how much interest will you receive in year 1 and year 2?
Answer #1

Face value of bond = $1,000
tenure of bond = 20 years
For equal annual amortization,
Principal to be received each year = $1,000/20 = $50

If the coupon rate is 2.8%/year
interest for year 1 = $1,000 *2.8% = $28
Interest for year 2 = ($1,000 - $50) * 2.8% = $26.6